OpenStudy (anonymous):

What is the value of the expression when a=6 and b=4? 14a-2(3b+4)

OpenStudy (marcelie):

You just plug in the variables for the equation

OpenStudy (anonymous):

a=6 and b=4? 14a-2(3b+4) Step 1: (3x4=12) step2:(12+4)=16 Step3:14x6=84 Step4:84-2=82 Step5:82+16=98 so i think the answer is 98

OpenStudy (anonymous):

@Dolarom

OpenStudy (anonymous):

@Cookie_2046 do you agree

OpenStudy (cookie_2046):

yes i agree (:

OpenStudy (wolf1728):

a=6 and b=4? 14a-2(3b+4) 14*6 -2*(3*4 +4) 84 - 2 * (16) 84 -32 = 52 I thought parentheses always take precedence

OpenStudy (wolf1728):

Do you agree cookie?

OpenStudy (cookie_2046):

yes i agree (:

OpenStudy (wolf1728):

So it is 52? Somebody better tell Dolarom

OpenStudy (jmark):

The value is 48.